The fixed costs of operating a business are the costs incurred regardless of the level of production. Fixed costs include rent, fixed salaries, and costs of leasing machinery.The variable costs of operating a business are the costs that change with the level of output. Variable costs include raw materials, hourly wages,and electricity. Suppose that a manufacturer of jeans has fixed daily costs of $500 and variable costs of $8 for each pair of jeans manufactured. Write a linear equation that relates the daily cost C, in dollars, of manufacturing the jeans to the number x of jeans manufactured. What is the cost of manufacturing 400 pairs of jeans? 740 pairs?
Found 2 solutions by ikleyn, mananth:
Answer by ikleyn(52776)

.
The fixed costs of operating a business are the costs incurred regardless of the level of production.
Fixed costs include rent, fixed salaries, and costs of leasing machinery.
The variable costs of operating a business are the costs that change with the level of output.
Variable costs include raw materials, hourly wages, and electricity.
Suppose that a manufacturer of jeans has fixed daily costs of $500 and variable costs of $8
for each pair of jeans manufactured.
(a) Write a linear equation that relates the daily cost C, in dollars, of manufacturing the jeans
to the number x of jeans manufactured.
(b) What is the cost of manufacturing 400 pairs of jeans? 740 pairs?
~~~~~~~~~~~~~~~

(a)  They provide you a nice description of all relevant notions and quantities in finance and economy.

     So, according to this description, your linear equation is

           C = 8x + 500  dollars daily.

     
     The formula says that the daily cost rises from 500 dollars by 8 dollars with each manufactured pair of jeans.



(b)  To calculate the daily cost at 400 pair of jeans, substitute x= 400 into the formula and get

            C(400) = 8*400 + 500 = 3700  dollars.


     Now make the same for 740 pairs.  You may report your answer to me as your comment through the TY window.


If you do, then refer to the problem ID, which is 1207852.


Otherwise, I will not know from whom I have this comment/message.

Solved.

-----------------

Happy learning !

Come again to this forum soon to learn something new.



Answer by mananth(16946)

a manufacturer of jeans has fixed daily costs of $500 and variable costs of $8 for each pair of jeans manufactured. Write a linear equation that relates the daily cost C, in dollars, of manufacturing the jeans to the number x of jeans manufactured. What is the cost of manufacturing 400 pairs of jeans? 740 pairs?
The fixed daily costs are same whether you manufacture or not.
fixed daily costs of $500
x jeans manufactured
The variable cost is related to numbesr produced
C= 500 +8x
You can do the rest
In first qquestion x= 400
in second 750
Plug in for x you have the cost